K. Baggett, C. Hovater, R. Geng, A. Hutton, G. Krafft, A. Kimber, K. Macha, C. Mounts, R. Nelson, T. Plawski, T. Powers, R. Rimmer, C. Reece, A. Seryi, and H. Wang.
Agenda and Notes
1. Operations Status and Summary/Review (Bickley)
Physics run extended and two weeks added to shutdown. Good reliability running at low energy. Re-planned summer maintenance work. Full E' plan will be executed. Low energy setup with reduced heating accomplished.
Plan for module swaps complete.
2. In Situ Q_L Measurements (Marhauser)
3. Some Comments on particulates (Geng)
Many valve closure tests completed and log entries made. Particulates ranging in size from sub-micron to several microns are observed in the tests.
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?